A bus starts from rest at a station and accelerates at a speed rate of 2m/s^2 for 15 seconds. It then runs at a constant speed for 25 seconds and slows down at -3m/s^2 until it stops. Find the total displacement travelled by the bus

1
Expert's answer
2021-11-07T17:06:52-0500

1) Displacement during acceleration:


"d_1=a_1t_1^2/2=225\text{ m}."


2) Displacement during the 25 s:


"d_2=vt_2=(a_1t_1)t_2=750\text{ m}."

3) Displacement during deceleration:


"d_3=v^2/(2a_3)=(a_1t_1)^2/(2a_3)=150\text{ m}."


4) The total displacement:


"D=d_1+d_2+d_3=1125\text{ m}."